A virus reproduces by dividing into two, and after a certain growth period, it dividing into two again. As the virus continue to divine in two. How many viruses will be in system starting with a single virus AFTER 10 division?
Answer by ikleyn(53763)

.
A virus reproduces by dividing into two, and after a certain growth period, it dividing into two again.
As the virus continue to divine in two. How many viruses will be in system starting with a single virus AFTER 10 division?
~~~~~~~~~~~~


After 1st division, there are 2 viruses.


After 2nd division, there are 2*2 = 2^2 = 4 = 2^2 viruses.


After 3rd division, there are 2*2^2 = 2*4 = 8 = 2^3 viruses.


After 4th division, there are 2*2^3 = 2*8 = 16 = 2^4 viruses.


   . . . . and so on . . . . 


After 10th division, there are 2*2^9 = 2*512 = 1024 = 2^10 viruses.      ANSWER

Solved.